# Novel insight into centre-vortex geometry in four dimensions

### v1

@@ I Introduction
One aspect of vacuum structure that has been well established as a leading prospect for a confinement mechanism is the presence of centre vortices in the ground-state fields [4, 5, 6, 7, 8]. These form closed two-dimensional sheets in four dimensions, and have been studied extensively in the framework of $\mathrm{SU}(N)$ lattice gauge theory [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. The removal of centre vortices has been demonstrated to result in a vanishing string tension, indicating the loss of an area law [25, 36]. In addition, centre-vortex removal with dynamical fermions leaves no sign of positivity violation in the Scwhinger function [37]. These findings highlight the importance of centre vortices in understanding the nonperturbative phenomena of QCD.
@@
Figure 1: Adapted from Ref. [40]. An illustration of how disconnected loops can arise by slicing through a single curved surface. Here, the $\tau$ coordinate is being held fixed such that we are looking at an $xy$ cross section. Depending on the slice coordinate, we can see either a single closed curve (top) or two disconnected curves (bottom).
@@ III Centre-vortex structures in four dimensions
Figure 2: Illustration for identifying connecting plaquettes in three-dimensional slices of the lattice. The middle $ij$ plaquette at $\mathbf{x}$ is pierced by an $m=+1$ vortex, rendered as a unit arrow on the dual lattice pointing in the $+\mathbf{\hat{k}}$ direction and piercing the plaquette. Due to conservation of centre charge, there must be at least one vortex connection to the left, and at least one to the right. Consequently, one can exclusively check the faces of the elementary cubes at $\mathbf{x}$ and $\mathbf{x-\hat{k}}$ to identify the next nontrivial plaquettes. This holds also if $m=-1$ and the vortex were to point in the $-\mathbf{\hat{k}}$ direction. In four dimensions, one need simply repeat this procedure for the two dimensions orthogonal to the pierced plaquette.
@@
Moving to four dimensions, there are now two orthogonal directions $\kappa$, $\lambda$ for a given $\mu\nu$ plaquette ($\epsilon_{\mu\nu\kappa\lambda}=1$). Noting that the arrows plotted on the dual lattice (as in Fig. 2) lie within the sheet, here the connecting nontrivial plaquettes can be identified by implementing the above three-dimensional procedure independently for $\kappa$ and $\lambda$. First, $\lambda$ is held fixed and one imagines the elementary cubes formed by translating the $\mu\nu$ plaquette one step forward and backward in the $\kappa$ direction. The plaquettes that form these cubes are checked for vortices. Then, the roles of $\kappa$ and $\lambda$ are interchanged—$\kappa$ is held fixed, and one checks the forward and backward elementary cubes in the $\lambda$ direction. After the faces of all four elementary cubes have been visited, all possible connecting plaquettes have been accounted for. This can then be applied recursively to map out centre-vortex sheets in four dimensions.
@@
The visualisations are constructed following Ref. [35]. We slice through a dimension of the four-dimensional lattice by holding the selected coordinate fixed. Slices obtained by fixing temporal and spatial coordinates are referred to as “temporal” and “spatial” slices, respectively. The orientation of an $m=+1$ vortex is determined by applying the right-hand rule, with an $m=-1$ vortex rendered in the opposite direction. Since the flow of $m=-1$ centre charge is indistinguishable from an opposite flow of $m=+1$ centre charge, this implies the visualisations exclusively show the flow of $m=+1$ centre charge. This convention is demonstrated in Fig. 3.
Figure 3: The visualisation convention for centre vortices. An $m=+1$ vortex (left) is represented by a jet in the available orthogonal dimension, with the direction given by the right-hand rule. An $m=-1$ vortex (right) is rendered by a jet in the opposite direction, such that we always show the flow of $m=+1$ centre charge.
@@ IV.1 Secondary sheets
The next-smallest sheets comprise ten vortices. These primarily constitute (but are not limited to) $1\times 1\times 2$ closed surfaces, and permutations thereof. Beyond a size of ten, there is an alternating trend where secondary sheets that comprise an even number of vortices are more probable than those with an odd number of vortices. This is because for a closed surface to contain an odd number of vortices, it necessarily features a branching point somewhere in the sheet. Centre-vortex branching points, also known as monopoles, are allowed because the centre charge is only conserved modulo $N$ in general $\mathrm{SU}(N)$ gauge theory. In $\mathrm{SU}(3)$, this implies that three $m=+1$ vortices can converge to or emerge from a point. The semantics of branching points are highlighted in Fig. 9.
Figure 9: Schematic of a vortex branching point/monopole. In four dimensions (left), monopoles occur when three faces of the vortex sheet connect at a common dual link. The black arrows around each face of the sheet are to indicate the vortex orientation, here taken to be $m=+1$. In taking a cross section through the sheet (i.e. “slicing”), as portrayed, the monopole appears as three vortex lines emerging from a point (right). These arrows appear where the cross section intersects the vortex sheet. Examples of such monopoles can be found throughout the visualisations in Figs. 4 and 5. Reversal of the left-hand arrow in the right figure instead shows the flow of $m=-1$ centre charge, which is in turn equivalent to $m=+2$ since $-1=+2\pmod{3}$. Therefore, the right-hand diagram equivalently depicts the branching of centre charge.
They are relatively rare occurrences, such that sheets with an odd number of vortices are suppressed.
@@ IV.2 Connected and disconnected loops
In contrast, the typical number of disconnected secondary loop above $T_{d}$ is roughly unchanged from below $T_{d}$, still only $\simeq 1$ per slice. As with previous quantities, this is excluding our temperature just above $T_{d}$. In fact, the trend displayed by the number of disconnected secondary loops follows directly from the secondary-sheet density in Fig. 7. That is, if there is a greater density of secondary sheets in four dimensions, then there will on average be a greater number of disconnected secondary loops in three-dimensional slices. One can even derive an approximate relation between the two quantities. Denoting the secondary-sheet density by $\rho_{\mathrm{sheet}}$, the number of secondary sheets is
@@
Furthermore, Fig. 12 reveals additional context for the anomaly barely above $T_{d}$. A similar increase is in fact also present below $T_{d}$, but in the connected secondary loops. It is now apparent that as $T_{d}$ is approached from below, there is additional curvature in the primary sheet before the alignment sets in that underlies this increase. Then, immediately as the phase transition is crossed, these additional regions of curvature effectively “break off” from the primary sheet, generating a temporary spike in secondary sheets. This in turn realises an increase in the number of disconnected secondary loops found in temporal slices. The fact that this dynamic is only found in a small region around $T_{d}$ suggests it would be interesting to investigate the nature of the finite-volume crossover in greater detail. We find it fascinating that our previous observations around $T_{d}$ can be understood as curvature “separating” from the primary sheet.
@@
Figure 15: Same as Fig. 14, but excluding the elementary $1\times 1$ vortex loops of length four. The asymmetry between even and odd lengths is apparent, with the latter suppressed due to requiring vortex branching. Note the differing vertical axis scales.

### v2

@@ I Introduction
One aspect of vacuum structure that has been well established as a leading prospect for a confinement mechanism is the presence of centre vortices in the ground-state fields [4, 5, 6, 7, 8]. These form closed two-dimensional sheets in four dimensions, and have been studied extensively in the framework of $\mathrm{SU}(N)$ lattice gauge theory [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. The removal of centre vortices has been demonstrated to result in a vanishing string tension, indicating the loss of an area law [17, 25, 36]. In addition, centre-vortex removal with dynamical fermions leaves no sign of positivity violation in the Schwinger function [37]. These findings highlight the importance of centre vortices in understanding the nonperturbative phenomena of QCD.
@@
Figure 1: Adapted from Ref. [40]. An illustration of how disconnected loops can arise by slicing through a single curved surface. Here, the $\tau$ coordinate is being held fixed such that we are looking at an xyxy cross section. Depending on the slice coordinate, we can see either a single closed curve (top) or two disconnected curves (bottom).
@@ III Centre-vortex structures in four dimensions
Figure 2: Illustration for identifying connecting plaquettes in three-dimensional slices of the lattice. The middle ijij plaquette at $\mathbf{x}$ is pierced by an $m=+1$ vortex, rendered as a unit arrow on the dual lattice pointing in the $\mathbf{\hat{k}}$ direction and piercing the plaquette. Due to conservation of centre charge, there must be at least one vortex connection to the left, and at least one to the right. Consequently, one can exclusively check the faces of the elementary cubes at $\mathbf{x}$ and $\mathbf{x-\hat{k}}$ to identify the next nontrivial plaquettes. This holds also if m=1m=-1 and the vortex were to point in the $\mathbf{\hat{k}}$ direction. In four dimensions, one need simply repeat this procedure for the two dimensions orthogonal to the pierced plaquette.
@@
[…]
@@
The visualisations are constructed following Ref. [35]. We slice through a dimension of the four-dimensional lattice by holding the selected coordinate fixed. Slices obtained by fixing temporal and spatial coordinates are referred to as “temporal” and “spatial” slices, respectively. The orientation of an $m=+1$ vortex is determined by applying the right-hand rule, with an m=1m=-1 vortex rendered in the opposite direction. Since the flow of m=1m=-1 centre charge is indistinguishable from an opposite flow of $m=+1$ centre charge, this implies the visualisations exclusively show the flow of $m=+1$ centre charge. This convention is demonstrated in Fig. 3.
Figure 3: The visualisation convention for centre vortices. An $m=+1$ vortex (left) is represented by a jet in the available orthogonal dimension, with the direction given by the right-hand rule. An m=1m=-1 vortex (right) is rendered by a jet in the opposite direction, such that we always show the flow of $m=+1$ centre charge.
@@ IV.1 Secondary sheets
Figure 9: Schematic of a vortex branching point/monopole. In four dimensions (left), monopoles occur when three faces of the vortex sheet (purple areas) connect at a common dual link. The black arrows around each face of the sheet are to indicate the vortex orientation, taken to be $m=+1$. In taking a cross section through the sheet (i.e. “slicing”), as portrayed by the dashed grey area, the monopole appears as three vortex lines emerging from a point (right). These arrows appear where the cross section intersects the vortex sheet. Examples of such monopoles can be found throughout the visualisations in Figs. 4 and 5. Reversal of the left-hand arrow in the right figure instead shows the flow of m=1m=-1 centre charge, which is in turn equivalent to $m=+2$ since $1=+2\pmod{3}$. Therefore, the right-hand diagram equivalently depicts the branching of centre charge.
The next-smallest sheets comprise ten vortices. These primarily constitute (but are not limited to) $1\times 1\times 2$ closed surfaces, and permutations thereof. Beyond a size of ten, there is an alternating trend where secondary sheets that comprise an even number of vortices are more probable than those with an odd number of vortices. This is because for a closed surface to contain an odd number of vortices, it necessarily features a branching point somewhere in the sheet. Centre-vortex branching points, also known as monopoles, are allowed because the centre charge is only conserved modulo $N$ in general $\mathrm{SU}(N)$ gauge theory. In $\mathrm{SU}(3)$, this implies that three $m=+1$ vortices can converge to or emerge from a point. The semantics of branching points are highlighted in Fig. 9. They are relatively rare occurrences, such that sheets with an odd number of vortices are suppressed.
@@ IV.2 Connected and disconnected loops
In contrast, the typical number of disconnected secondary loop above $T_{d}$ is roughly unchanged from below $T_{d}$, still only $\simeq 1$ per slice. As with previous quantities, this is excluding our temperature just above $T_{d}$, which is slightly higher. In fact, the trend displayed by the number of disconnected secondary loops follows directly from the secondary-sheet density in Fig. 7. That is, if there is a greater density of secondary sheets in four dimensions, then there will on average be a greater number of disconnected secondary loops in three-dimensional slices. One can even derive an approximate relation between the two quantities. Denoting the secondary-sheet density by $\rho_{\mathrm{sheet}}$, the number of secondary sheets is
@@
Furthermore, Fig. 12 reveals additional context for the anomaly barely above $T_{d}$. A similar increase is in fact also present below $T_{d}$, but in the connected secondary loops. We also know from Ref. [39] that there is a slight increase in vortex density just below $T_{d}$. Together, this makes it apparent that as $T_{d}$ is approached from below, the extra vortex matter manifests as additional curvature in the primary sheet. This underlies the increase in connected secondary loops in this region below $T_{d}$. Then, immediately as the phase transition is crossed and the alignment sets in, these additional regions of curvature effectively “break off” from the primary sheet, generating a temporary spike in secondary sheets. This in turn realises an increase in the number of disconnected secondary loops found in temporal slices. The fact that this dynamic is only found in a small region around $T_{d}$ suggests it would be interesting to investigate the finite-volume crossover in greater detail. We find it fascinating that our previous observations around $T_{d}$ can be understood as curvature “separating” from the primary sheet.
@@
Figure 15: Same as in Fig. 14, but excluding the elementary $1\times 1$ vortex loops of length four. The asymmetry between even and odd lengths is apparent at all temperatures, with the later suppressed due to requiring vortex branching.